Constants Periodic Table Part A A 32.0 g iron rod, initially at 21.8 °C, is submerged into an unknown mass of water at 63.5° C, in an insulated container. The final temperature of the mixture upon reaching thermal equilibrium is 59.6 C. What is the mass of the water? Express your answer to two significant figures and include the appropriate units You may want to reference (Pages 259 - 264) Section 6.4 while completing this problem m Value Units
Answer #1

let 1 denote Fe and 2 denote water
s1 = 0.452 J/g.oC
s2 = 4.186 J/g.oC

use,
m1*s1*(T-T1) = m2*s2*(T2-T)
32.0*0.452*(59.6 - 21.8) = m2*4.186*(63.5 - 59.6)
32.0*0.452*37.8= m2*4.186*3.9
546.7 = m2*16.32
m2 = 33.5 g

Answer : 33.5 g